### v1

# backref=true

Abstract.
We prove that the derived categories of noncommutative curves in the sense of Reiten-van den Bergh do not contain phantom subcategories. This adds many examples not presently in the literature of derived categories without phantom subcategories.
1. Introduction
In algebraic geometry and representation theory, triangulated cat-egories are by now widely recognized as important tools. Originally defined as book-keeping devices, it was later realized that they are interesting objects in and of themselves [Beilinson1978] [BGG78] [MukaiAbelian]. Subseqently, there was an explosion of interest in the 90’s related to Kontsevich’s homological mirror symmetry proposal [kontsevich1994homological].
Semiorthogonal decompositions (SODs) are a way to understand triangulated categories by breaking them into smaller pieces. When one considers the bounded derived category of coherent sheaves of a variety, $\operatorname{D^{b}_{coh}}(X)$, these decompositions also often correlate to interesting geometric phenomena. As a consequence, conjectures were posed as to the possible structures of their basic building blocks, admissible sub-categories (2.1).
Definition 1.1.
For a triangulated category $\mathcal{D}$, an admissible sub-category $\mathcal{B}\subset\mathcal{D}$ is called a phantom if $K_{0}(\mathcal{B})=0$.
It was conjectured that phantom categories should not exist in non-pathological categories. However, in the last decade examples have been found in bounded derived categories of smooth projective surfaces. Quite recently, Krah found an example on a rational surface [krah2023phantom]. Such a surface has a tilting bundle, which means that this example also gives an example of a finite dimensional algebra $A$ such that $\mathrm{D}^{\mathrm{b}}(\mathrm{mod}\>A)$ has a phantom, as observed in [kalck2023finite].
Given these results, one might begin to view existence of phantoms as the rule rather than the exception. Indeed, to date very few non-existence results have been proven. It is folklore that for a smooth projective curve $X$, $\operatorname{D^{b}_{coh}}(X)$ contains no phantom subcategories. Pirozhkov [pirozhkov2020admissible] proves that for a del Pezzo surface $S$, $\operatorname{D^{b}_{coh}}(S)$ does not contain a phantom. In [PirozhkovstablySO] several other classes of examples are ruled out including products of the form $C\times\mathbb{P}^{1}$ with $C$ a positive genus curve.
For an Abelian category $\mathcal{A}$, its homological dimension is
$\mathrm{hd}(\mathcal{A})=\sup_{E,F\in\mathcal{A}}\{n:\operatorname{Ext}^{n}_{\mathcal{A}}(E,F)\neq 0\}.$
Here, $\operatorname{Ext}^{n}_{\mathcal{A}}(E,F):=\operatorname{Hom}_{\mathrm{D}(\mathcal{A})}(E,F[n]).$ This agrees with the usual definition when $\mathcal{A}$ has enough injectives or projectives.
In the present note, we prove that for many Abelian categories $\mathcal{A}$ of homological dimension $1$, $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ does not contain a phantom sub-category. In particular, all noncommutative curves in the sense of [ReitenvDB] do not contain phantom subcategories. This result is probably known to the experts, however it does not seem to appear anywhere in the literature.
Acknowledgements:
I thank Matthew Ballard for first encouraging me to work out the case of an acylic quiver and Pieter Belmans for useful conversations and for his interest in this project. I am also indebted to Daniel Krashen, Svetlana Makarova, Alexander Perry, and the rest of my group at the AMS MRC program for stimulating discussions about this topic. I thank Yuri Berest for his helpful input concerning this project. I am especially grateful to Andres Fernandez Herrero for his continued interest and helpful suggestions; it was his suggestion to consult [Nironi09]. Finally, it is my pleasure to thank my advisor Daniel Halpern-Leistner for his continued guidance and support.
2. The main lemma
Definition 2.1.
[B-KSerre] Let $\mathcal{D}$ be a triangulated category and consider a full triangulated subcategory $\mathcal{B}$ of $\mathcal{D}$.
  • $\mathcal{B}^{\perp}$ denotes the full subcategory of $\mathcal{D}$ consisting of objects $X$ such that $\operatorname{Hom}_{\mathcal{D}}(B,X)=0$ for all $B\in\mathcal{B}$
  • ${}^{\perp}\mathcal{B}$ denotes the full subcategory of $\mathcal{D}$ consisting of objects $X$ such that $\operatorname{Hom}_{\mathcal{D}}(X,B)=0$ for all $B\in\mathcal{B}$.
We say that $\mathcal{B}$ is
  1. (1)
    left admissible if every $X\in\mathcal{D}$ fits into a distinguished triangle $A\to X\to B$ with $A\in{}^{\perp}\mathcal{B}$ and $B\in\mathcal{B}$;
  2. (2)
    right admissible if every $X\in\mathcal{D}$ fits into a distinguished triangle $B\to X\to A$ with $B\in\mathcal{B}$ and $A\in\mathcal{B}^{\perp}$; and
  3. (3)
    admissible if it is both left and right admissible.
Lemma 2.2.
If $\mathcal{B}$ is an admissible subcategory of $\mathcal{D}$ then $X\oplus Y\in\mathcal{B}$ implies $X\in\mathcal{B}$.
Proof.
This follows from $\mathcal{B}={}^{\perp}(\mathcal{B}^{\perp})$; see [stacks-project, Section 0CQP]. ∎
For an Abelian category $\mathcal{A}$, its bounded derived category $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ has a standard bounded t-structure with t-cohomology functors denoted $\{H^{i}\}_{i\in\mathbb{Z}}$. In particular, for any $X\in\mathrm{D}^{\mathrm{b}}(\mathcal{A})$, $H^{i}(X)\in\mathcal{A}$ for all $i$.
Lemma 2.3.
If $\mathcal{A}$ is an Abelian category of homological dimension $1$, then for every object $X\in\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ one has $X\cong\bigoplus_{i\in\mathbb{Z}}H^{i}(X)[-i]$, where all but finitely many summands are zero.
Proof.
See [HuybrechtsFM]*Cor. 3.15. ∎
Definition 2.4.
An Abelian category $\mathcal{A}$ is called visible if for all $X\in\mathcal{A}$, $[X]=0\in K_{0}(\mathcal{A})$ implies $X=0$.
The following simple lemma is the fundamental observation.
Lemma 2.5.
If $\mathcal{A}$ is a homological dimension $1$ visible Abelian category then $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ contains no phantom subcategory.
Proof.
Suppose $\mathcal{B}\subset\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ is an admissible subcategory. Let $X\in\mathcal{B}$ be given and decomposed as in 2.3. If $X\neq 0$, then there exists an $i$ such that $H^{i}(X)\neq 0$ and by 2.2 one has $\mathcal{B}\cap\mathcal{A}\neq 0$. If $\mathcal{B}$ were a phantom then its objects would have zero image in $K_{0}(\mathcal{A})$. In particular, any $X\in\mathcal{B}\cap\mathcal{A}$ must have trivial class in $K_{0}(\mathcal{A})$. By hypothesis, this implies $\mathcal{B}\cap\mathcal{A}=0$ and hence that $\mathcal{B}=0$. ∎
3. Examples
We next prove that examples of interest satisfy the visible condition.
3.1. Hereditary finite dimensional algebras
An Abelian category $\mathcal{A}$ is called a Jordan-Hölder (JH) category if every $A\in\mathcal{A}$ is Noetherian and Artinian. It follows that each $A\in\mathcal{A}$ has a filtration $0=A_{0}\subset A_{1}\subset\cdots\subset A_{n}=A$ such that
  1. (1)
    the associated graded objects $S_{i}=A_{i}/A_{i-1}$ for $1\leq i\leq n$ are simple; and
  2. (2)
    the set of isomorphism classes of the associated graded objects is independent of the filtration.
For proofs, see [stacks-project, Tag 0FCJ].
Lemma 3.1.
If $\mathcal{A}$ is a JH Abelian category then $\mathcal{A}$ is visible.
Proof.
$K_{0}(\mathcal{A})=\bigoplus_{S\>\text{simple}}\mathbb{Z}\cdot[S]$ and given $M\in\mathcal{A}$, $[M]=\sum_{i=1}^{k}n_{i}[S_{i}]$ where $S_{1},\ldots,S_{k}$ are the simples appearing in any composition series of $M$ with multiplicities $n_{1},\ldots,n_{k}$, respectively. Visibility follows. ∎
For $A$ an algebra over a field $k$, $\mathrm{mod}\>A$ denotes the category of finite dimensional representations of $A$.
Corollary 3.2.
Let $A$ denote a hereditary finite dimensional algebra over a field $k$. $\mathrm{D}^{\mathrm{b}}(\mathrm{mod}\>A)$ contains no phantom subcategory.
Proof.
$A$ hereditary means $\mathrm{hd}(\mathrm{mod}\>A)=1$. $\mathrm{mod}\>A$ is a Jordan-Hölder category so we apply 3.1. ∎
Example 3.3.
Let $Q$ denote an acyclic quiver and $\mathrm{rep}_{k}(Q)$ its category of finite dimensional $k$-representations. $\mathrm{rep}_{k}(Q)$ is equivalent to the category of finite dimensional modules over the path algebra $kQ$, a finite dimensional hereditary algebra over $k$. So, $\mathrm{D}^{\mathrm{b}}(\mathrm{rep}_{k}(Q))$ contains no phantoms by 3.2.
Example 3.4.
Let $X$ denote a smooth and proper curve over $\operatorname{Spec}k$. $\operatorname{Coh}(X)$ is visible because for $\mathcal{F}\in\operatorname{Coh}(X)$, $\operatorname{rank}(\mathcal{F})=\deg(\mathcal{F})=0$ implies $\mathcal{F}=0$.
Remark 3.5.
It is folklore that $\operatorname{D^{b}_{coh}}(X)$ does not contain phantoms for $X$ a smooth and proper curve. However, the typical argument is less intrinsic. For $\mathbb{P}^{1}$, one can apply the classification of SODs of $\operatorname{Coh}(\mathbb{P}^{1})$. For $g(X)=1$, $\operatorname{D^{b}_{coh}}(X)$ is a Calabi-Yau category and hence indecomposable (i.e. contains no admissible subcategories). For $g(X)\geq 2$, Okawa’s theorem [Okawaindecomp] implies $\operatorname{D^{b}_{coh}}(X)$ is indecomposable.
3.2. Stacky curves
We suppose throughout this section and the next that $k$ is algebraically closed and of characteristic zero.
Definition 3.6.
A stacky curve $\mathcal{X}$ is an irreducible one dimensional Deligne-Mumford (DM) stack over $k$ with generically trivial stabilizers.
Definition 3.7.
A DM stack $\mathcal{X}$ is called projective if it admits a closed embedding into a smooth proper DM stack with projective coarse moduli space. (See [Kresch]*Thm. 5.3.)
For a smooth projective stacky curve $\mathcal{X}$, it is not hard to see that its category of coherent sheaves, $\operatorname{Coh}(\mathcal{X})$, is of homological dimension $1$. Indeed, it follows from [NironiDuality]*Thm 2.22 that $\operatorname{D^{b}_{coh}}(\mathcal{X})$ has a Serre functor given by $\omega_{\mathcal{X}}[1]$, where $\omega_{\mathcal{X}}$ is the canonical sheaf of $\mathcal{X}$. Then for all $\mathcal{F},\mathcal{G}\in\operatorname{Coh}(\mathcal{X})$ one has
$\operatorname{Ext}^{i}_{\mathcal{O}_{\mathcal{X}}}(\mathcal{F},\mathcal{G})\cong\operatorname{Ext}^{1-i}_{\mathcal{O}_{\mathcal{X}}}(\mathcal{G},\mathcal{F}\otimes\omega_{\mathcal{X}})^{*}=0\>\>\text{for all}\>\>i\geq 2.$
One can also show that the coarse moduli space of a projective DM stack is projective [Kresch]*Cor 5.4.
Proposition 3.8.
If $\mathcal{X}$ is a projective DM stack, then there exists a group homomorphism $P:K_{0}(\mathcal{X})\to\mathbb{Q}[t]$ such that for $\mathcal{F}\in\operatorname{Coh}(\mathcal{X})$, $P(\mathcal{F})=0$ implies $\mathcal{F}=0$.
Proof.
Let $\pi:\mathcal{X}\to X$ denote the coarse moduli space of $\mathcal{X}$. For a locally free generating sheaf $\mathcal{E}$ on $\mathcal{X}$ (such a sheaf always exists by the projectivity assumption [Kresch]*Thm. 5.3), there is a functor $F_{\mathcal{E}}:\operatorname{QCoh}(\mathcal{X})\to\operatorname{QCoh}(X)$ given by $\mathcal{F}\mapsto\pi_{*}\operatorname{Hom}_{\mathcal{O}_{\mathcal{X}}}(\mathcal{E},\mathcal{F})$ [Nironi09]*Defn. 2.4. As $\pi$ is proper, $F_{\mathcal{E}}$ induces a functor $\operatorname{Coh}(\mathcal{X})\to\operatorname{Coh}(X)$.
$F_{\mathcal{E}}$ is exact and for all $\mathcal{F}\in\operatorname{Coh}(\mathcal{X})$, $F_{\mathcal{E}}(\mathcal{F})=0$ if and only if $\mathcal{F}=0$ [Nironi09]*Lem. 3.4. Fix a very ample line bundle $\mathcal{O}_{X}(1)$ on $X$. For $\mathcal{G}\in\operatorname{Coh}(X)$ let $\mathcal{G}(m):=\mathcal{G}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{X}(1)^{\otimes m}$. Using this, one defines $P_{\mathcal{E}}(\mathcal{F},m):=\chi(X,F_{\mathcal{E}}(\mathcal{F})(m))$, which is a rational polynomial function of $m\in\mathbb{Z}$ and defines a polynomial $P_{\mathcal{E}}(\mathcal{F},t)\in\mathbb{Q}[t]$. This gives a homomorphism $P_{\mathcal{E}}(-):K_{0}(\mathcal{X})\to\mathbb{Q}[t]$. Now, $P_{\mathcal{E}}(\mathcal{F},t)\in\mathbb{Q}[t]$ is the Hilbert polynomial of $F_{\mathcal{E}}(\mathcal{F})\in\operatorname{Coh}(X)$; so, if $P_{\mathcal{E}}(\mathcal{F},t)=0$, then $F_{\mathcal{E}}(\mathcal{F})=0$ and hence $\mathcal{F}=0$. ∎
Corollary 3.9.
If $\mathcal{X}$ is a projective stacky curve, then $\operatorname{Coh}(\mathcal{X})$ is visible. Consequently, $\operatorname{D^{b}_{coh}}(\mathcal{X})$ contains no phantom subcategories.
Proof.
$\operatorname{Coh}(\mathcal{X})$ is visible by 3.8; indeed, for $\mathcal{F}\in\operatorname{Coh}(\mathcal{X})$, if $[\mathcal{F}]=0$, then $P_{\mathcal{E}}(\mathcal{F})=0$ and $\mathcal{F}=0$. We conclude by 2.5. ∎
We conclude with a more specific example.
Example 3.10.
A weighted projective line $\mathcal{X}$ is a smooth stacky curve with coarse moduli space $\mathbb{P}^{1}$. $\operatorname{D^{b}_{coh}}(\mathcal{X})$ has been studied in many guises: $\mathcal{X}$ admits a tilting bundle $E$ inducing an equivalence $\operatorname{D^{b}_{coh}}(\mathcal{X})\to\mathrm{D}^{\mathrm{b}}(\mathrm{mod}\>A)$ for $A=\operatorname{End}(E)$ a finite dimensional $k$-algebra. The study of the representation theory of the resulting algebras was initiated in [GeigleLenzing] and a vast literature exists. This class also includes the weighted projective spaces $\mathbb{P}(a,b)$ where $a,b\geq 1$ have $\gcd(a,b)=1$.
4. Noncommutative curves
A $k$-linear category $\mathcal{D}$ is called finite-type if $\operatorname{Hom}_{\mathcal{D}}(X,Y)$ is a finite dimensional $k$-vector space for all $X,Y\in\mathcal{D}$.
Definition 4.1.
A finite-type $k$-linear triangulated category $\mathcal{D}$ is called saturated if every (covariant and contravariant) cohomological functor $H:\mathcal{D}\to\mathrm{mod}(k)$ is representable.
Saturatedness implies existence of a Serre functor on $\mathcal{D}$, which gives rise to a version of Serre duality [B-KSerre]. By the results ibid., saturated triangulated categories are a natural setting in which to study admissible subcategories.
Definition 4.2.
[ReitenvDB] A noncommutative curve over $k$ is a $k$-linear Abelian category $\mathcal{A}$ with $\mathrm{hd}(\mathcal{A})=1$ such that $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ is saturated.
In [ReitenvDB] noncommutative curves over a field $k$ are classified as belonging to the following two classes:
  1. (1)
    $\mathrm{mod}\>\Lambda$, for $\Lambda$ a hereditary finite dimensional algebra over $k$; or
  2. (2)
    $\operatorname{Coh}(\mathcal{O})$, where $\mathcal{O}$ is a hereditary order on a smooth proper curve $X$ over $k$.
However, [ChanIngallsCoordinateRings]*Cor. 7.8 states that for all hereditary orders $\mathcal{O}$ on a smooth proper curve $X$, $\operatorname{Coh}(\mathcal{O})\simeq\operatorname{Coh}(\mathcal{X})$, where $\mathcal{X}$ is a smooth and projective stacky curve associated to $\mathcal{O}$. So, noncommutative curves are either equivalent to $\mathrm{mod}\>\Lambda$ as in (1) or $\operatorname{Coh}(\mathcal{X})$ for $\mathcal{X}$ a smooth and projective stacky curve over $k$.
Theorem 4.3.
If $\mathcal{A}$ is a noncommutative curve, then $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ contains no phantom subcategory.
Proof.
By the discussion above, it suffices to apply 3.2 and 3.9. ∎