### v2

# backref=true

Abstract.
We study admissible subcategories of the derived categories of smooth noncommutative curves as classified by Reiten-van den Bergh. We prove phantom subcategories do not exist in these derived categories and we obtain a complete classification of the admissible subcategories of the bounded derived category of representations of an acyclic quiver.
1. Introduction
In algebraic geometry and representation theory, triangulated categories are by now widely recognized as fundamental objects. Originally defined as book-keeping devices, it was later realized that they are interesting objects in and of themselves [Beilinson1978, BGG78, MukaiAbelian]. Subseqently, there was an explosion of interest in the 90’s related to Kontsevich’s homological mirror symmetry proposal [kontsevich1994homological].
Semiorthogonal decompositions, as introduced in [B-KSerre], are a way to understand triangulated categories by breaking them into smaller pieces. When one considers the bounded derived category of coherent sheaves of a variety, these decompositions also often correlate to interesting geometric phenomena. As a consequence, conjectures were posed as to the possible structures of their basic building blocks, admissible subcategories (see 3.1).
Definition 1.1.
For a triangulated category $\mathcal{D}$, an admissible subcategory $\mathcal{B}\subset\mathcal{D}$ is called a phantom if $K_{0}(\mathcal{B})=0$.
It was conjectured that phantom categories should not exist in non-pathological categories. However, in the last decade, examples have been found in bounded derived categories of smooth projective surfaces. Quite recently, Krah found an example on a rational surface [krah2023phantom]. Such a surface has a tilting bundle, which means that this example also gives an example of a finite dimensional algebra $A$ such that $\mathrm{D}^{\mathrm{b}}(\mathrm{mod}\>A)$ has a phantom [kalck2023finite].
Given these results, one might begin to view existence of phantoms as the rule rather than the exception. Indeed, to date few non-existence results have been proven. It is a folklore result that for a smooth projective curve $X$, $\operatorname{D^{b}_{coh}}(X)$ contains no phantom subcategories. Pirozhkov has proven that for a del Pezzo surface $S$, $\operatorname{D^{b}_{coh}}(S)$ does not contain a phantom [pirozhkov2020admissible]. In [PirozhkovstablySO], several other classes of varieties are proven to not admit phantoms, including $C\times\mathbb{P}^{1}$ with $C$ a positive genus curve. It was shown in [BorisovKemboi] that for varieties $X$ obtained by blowing up $\mathbb{P}^{2}$ along finitely many points in very general position on a cubic, $\operatorname{D^{b}_{coh}}(X)$ contains no phantom subcategories.
For an Abelian category $\mathcal{A}$, its homological dimension is
$\mathrm{hd}(\mathcal{A})=\sup_{E,F\in\mathcal{A}}\{n:\operatorname{Ext}^{n}_{\mathcal{A}}(E,F)\neq 0\}.$
Here, $\operatorname{Ext}^{n}_{\mathcal{A}}(E,F):=\operatorname{Hom}_{\mathrm{D}(\mathcal{A})}(E,F[n]).$ This agrees with the usual definition when $\mathcal{A}$ has enough injectives or projectives.
In the present note, we prove that for many Abelian categories $\mathcal{A}$ of homological dimension $1$, $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ does not contain a phantom subcategory. More precisely, we prove that for $\mathcal{A}$ a (smooth) noncommutative curve as studied in [ReitenvDB], any indecomposable admissible subcategory of $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ is equivalent to the bounded derived category of a noncommutative curve or to $\operatorname{D^{b}_{coh}}(\mathrm{pt})$.
As an application of this result, we prove that for $Q$ a finite acyclic quiver, any admissible subcategory of $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q)$ is equivalent to $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q^{\prime})$ for $Q^{\prime}$ another finite acyclic quiver. In particular, all admissible subcategories of $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q)$ admit full exceptional collections.
Acknowledgements
I thank Matthew Ballard for first encouraging me to think about phantom categories in derived categories of quivers. I am also indebted to Kimoi Kemboi, Daniel Krashen, Svetlana Makarova, Alexander Perry, and the rest of my group at the AMS MRC program for stimulating discussions about this topic. I also thank Pieter Belmans, Yuri Berest, Michel van den Bergh, and Severin Barmeier for helpful input concerning this project. I am especially grateful to Andres Fernandez Herrero for his continued interest and helpful suggestions. Finally, it is my pleasure to thank my advisor Daniel Halpern-Leistner for his continued guidance and support.
2. Noncommutative curves
A $k$-linear category $\mathcal{D}$ is called finite-type if $\operatorname{Hom}_{\mathcal{D}}(X,Y)$ is a finite dimensional $k$-vector space for all $X,Y\in\mathcal{D}$.
Definition 2.1.
A finite-type $k$-linear triangulated category $\mathcal{D}$ is called saturated if every (covariant and contravariant) cohomological functor $H:\mathcal{D}\to\mathrm{mod}\>k$ is representable.
Saturatedness implies existence of a Serre functor on $\mathcal{D}$, which gives rise to a version of Serre duality [B-KSerre]. By the results ibid., saturated triangulated categories are a natural setting in which to study admissible subcategories.
Definition 2.2.
[ReitenvDB] A (smooth) noncommutative curve over $k$ is a $k$-linear Noetherian Abelian category $\mathcal{A}$ with $\mathrm{hd}(\mathcal{A})=1$ such that $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ is saturated.
Saturatedness imposes a type of categorical smoothness. In what follows, we assume all noncommutative curves are smooth in this sense.
Example 2.3.
Let $A$ denote a finite dimensional algebra over a field $k$ and write $\mathrm{mod}\>A$ for the Abelian category of its finite dimensional modules. If $A$ is hereditary, i.e. any submodule of a projective $A$-module is projective, then $\mathrm{hd}(\mathrm{mod}\>A)\leq 1$ since every module admits a two term projective resolution.
Furthermore, it is a classical result that given a hereditary finite dimensional algebra $A$ there exists a finite acyclic quiver such that $\operatorname{rep}Q\cong\mathrm{mod}\>A$ [Assem_Skowronski_Simson_2006]*Ch. VII, Thm. 1.7. Then, one can verify that if there are two adjacent vertices in $Q$, the corresponding simple objects admit a nontrivial extension. So, for $Q$ a connected finite acyclic quiver, $\operatorname{rep}Q$ is a noncommutative curve if and only if $Q$ has at least $2$ vertices. Otherwise, $\operatorname{rep}Q=\mathrm{mod}\>k$.
Example 2.4.
Let $\mathcal{X}$ denote an orbifold curve over $k$, by which we mean an irreducible one-dimensional Deligne-Mumford stack with generically trivial stabilizers. A Deligne-Mumford stack $\mathcal{X}$ is called projective if it admits a closed embedding into a smooth and proper Deligne-Mumford stack with projective coarse moduli space [Kresch]*Thm. 5.3.
Consider $\operatorname{Coh}(\mathcal{X})$, the category of coherent sheaves on $\mathcal{X}$ where $\mathcal{X}$ is a projective orbifold curve in the sense just defined. $\mathcal{X}$ is Noetherian since it can be written in the form $[Y/G]$ for $Y$ a quasi-projective variety and $G$ a reductive algebraic group acting linearly on it (see [Kresch]). By [NironiDuality]*Thm 2.22, $\operatorname{D^{b}_{coh}}(\mathcal{X})$ has a Serre functor given by $-\otimes_{\mathcal{O}_{\mathcal{X}}}\omega_{\mathcal{X}}[1]$, where $\omega_{\mathcal{X}}$ is the canonical sheaf of $\mathcal{X}$. Then for all $\mathcal{F},\mathcal{G}\in\operatorname{Coh}(\mathcal{X})$,
$\operatorname{Ext}^{i}_{\mathcal{O}_{\mathcal{X}}}(\mathcal{F},\mathcal{G})\cong\operatorname{Ext}^{1-i}_{\mathcal{O}_{\mathcal{X}}}(\mathcal{G},\mathcal{F}\otimes\omega_{\mathcal{X}})^{\vee}=0\>\>\text{for all}\>\>i\geq 2.$
In particular, $\mathrm{hd}\operatorname{Coh}(\mathcal{X})\leq 1$. On the other hand, $\operatorname{Ext}^{1}(\mathcal{O}_{\mathcal{X}},\omega_{\mathcal{X}})=\operatorname{Hom}(\mathcal{O}_{\mathcal{X}},\mathcal{O}_{\mathcal{X}})^{\vee}\neq 0$ so that $\mathrm{hd}\operatorname{Coh}(\mathcal{X})=1$. $\operatorname{D^{b}_{coh}}(\mathcal{X})$ is saturated by [BerghLuntsSchnurer] and standard arguments show that $\operatorname{Coh}(\mathcal{X})$ is Noetherian. Thus, $\mathcal{X}$ is a noncommutative curve.
In fact, 2.3 and 2.4 are all of the “indecomposable” examples of noncommutative curves. We call an Abelian category $\mathcal{A}$ connected if given a decomposition $\mathcal{A}=\mathcal{A}_{1}\oplus\mathcal{A}_{2}$ with $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ full Abelian subcategories of $\mathcal{A}$, one has $\mathcal{A}_{1}=0$ or $\mathcal{A}_{2}=0$.
Theorem 2.5.
[ReitenvDB, ChanIngallsCoordinateRings] Let $\mathcal{A}$ denote a connected Noetherian Abelian category with $\mathrm{hd}(\mathcal{A})\in\{0,1\}$ such that $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ is saturated. Then $\mathcal{A}$ is equivalent to
  1. (1)
    $\operatorname{rep}Q$ for $Q$ a finite acyclic quiver; or
  2. (2)
    $\operatorname{Coh}(\mathcal{X})$ for $\mathcal{X}$ a smooth projective orbifold curve.
Moreover, if $\mathrm{hd}(\mathcal{A})=0$ then $\mathcal{A}=\mathrm{mod}\>k$.
Proof.
By [ReitenvDB]*Thm. V.1.2, $\mathcal{A}$ is either equivalent to the category of finite dimensional modules over a hereditary algebra or to $\operatorname{Coh}(\mathcal{O}_{X})$, where $\mathcal{O}_{X}$ is a sheaf of hereditary orders over a nonsingular connected projective curve $X$. In the former case, one applies [Assem_Skowronski_Simson_2006]*Ch. VII, Thm. 1.7 to obtain (1). In the latter case, one applies the results of [ChanIngallsCoordinateRings] which state that $\operatorname{Coh}(\mathcal{O}_{X})$ is equivalent to the category of coherent sheaves on an associated projective smooth orbifold curve $\mathcal{X}$. Finally, suppose $\mathrm{hd}(\mathcal{A})=0$. In light of the discussions in 2.3 and 2.4, this is only possible when $\mathcal{A}$ is the category of finite dimensional representations of a quiver with one vertex and no arrows, i.e. if $\mathcal{A}=\mathrm{mod}\>k$ as claimed. ∎
We conclude this section by characterizing the categories of type (1) in 2.5.
Definition 2.6.
An Abelian category $\mathcal{A}$ is called finite length if each $A\in\mathcal{A}$ is Noetherian and Artinian. It follows that each $A\in\mathcal{A}$ has a filtration $0=A_{0}\subset A_{1}\subset\cdots\subset A_{n}=A$ such that
  1. (1)
    the associated graded objects $S_{i}=A_{i}/A_{i-1}$ for $1\leq i\leq n$ are simple; and
  2. (2)
    the set of isomorphism classes of the associated graded objects and their multiplicities are independent of the filtration.
For proofs, see [stacks-project, Tag 0FCJ].
Lemma 2.7.
If $\mathcal{A}$ is a noncommutative curve of finite length, then $\mathcal{A}$ is equivalent to the category of finite dimensional modules over a finite dimensional hereditary algebra.
Proof.
Let $A$ denote a finite dimensional $k$-algebra. One can check that $\mathrm{mod}\>A$ is a finite length Abelian category by arguing with dimension of submodules. By 2.5, it suffices to show that $\operatorname{Coh}(\mathcal{X})$ is never finite length for $\mathcal{X}$ an orbifold curve. For this, consider $x\in\mathcal{X}$ a closed point with trivial automorphism group for simplicity. There is an infinite descending chain
$\cdots\hookrightarrow\mathcal{O}_{\mathcal{X}}(-nx)\hookrightarrow\cdots\hookrightarrow\mathcal{O}_{\mathcal{X}}(-x)\hookrightarrow\mathcal{O}_{\mathcal{X}}$
of inclusions of sheaves, where $\mathcal{O}_{\mathcal{X}}(-x)\hookrightarrow\mathcal{O}_{\mathcal{X}}$ is the map arising from multiplying by a local equation cutting out $x$. $\mathrm{coker}(\mathcal{O}_{\mathcal{X}}(-x)\to\mathcal{O}_{\mathcal{X}})$ is a nonzero torsion sheaf and consequently this is an infinite descending chain of subobjects of $\mathcal{O}_{\mathcal{X}}$ which does not stabilize. ∎
3. The main theorem
Definition 3.1.
[B-KSerre] Let $\mathcal{D}$ be a triangulated category and consider a full triangulated subcategory $\mathcal{B}$ of $\mathcal{D}$.
  • $\mathcal{B}^{\perp}$ denotes the full subcategory of $\mathcal{D}$ consisting of objects $X$ such that $\operatorname{Hom}_{\mathcal{D}}(B,X)=0$ for all $B\in\mathcal{B}$
  • ${}^{\perp}\mathcal{B}$ denotes the full subcategory of $\mathcal{D}$ consisting of objects $X$ such that $\operatorname{Hom}_{\mathcal{D}}(X,B)=0$ for all $B\in\mathcal{B}$.
We say that $\mathcal{B}$ is
  1. (1)
    left admissible if every $X\in\mathcal{D}$ fits into a distinguished triangle $A\to X\to B$ with $A\in{}^{\perp}\mathcal{B}$ and $B\in\mathcal{B}$;
  2. (2)
    right admissible if every $X\in\mathcal{D}$ fits into a distinguished triangle $B\to X\to A$ with $B\in\mathcal{B}$ and $A\in\mathcal{B}^{\perp}$; and
  3. (3)
    admissible if it is both left and right admissible.
Lemma 3.2.
If $\mathcal{B}$ is an admissible subcategory of $\mathcal{D}$ then $X\oplus Y\in\mathcal{B}$ implies $X\in\mathcal{B}$.
Proof.
This follows from $\mathcal{B}={}^{\perp}(\mathcal{B}^{\perp})$; see [stacks-project, Section 0CQP]. ∎
For an Abelian category $\mathcal{A}$, its bounded derived category $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ has a standard bounded t-structure with t-cohomology functors denoted $\{H^{i}\}_{i\in\mathbb{Z}}$. In particular, for any $X\in\mathrm{D}^{\mathrm{b}}(\mathcal{A})$, $H^{i}(X)\in\mathcal{A}$ for all $i$. We always regard $\mathcal{A}$ as the full subcategory of $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ whose objects have cohomology concentrated in degree 00.
Lemma 3.3.
If $\mathcal{A}$ is an Abelian category with $\mathrm{hd}(\mathcal{A})\in\{0,1\}$, then for every object $X\in\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ one has $X\cong\bigoplus_{i\in\mathbb{Z}}H^{i}(X)[-i]$, where all but finitely many summands are zero.
Proof.
See [HuybrechtsFM]*Cor. 3.15. ∎
Lemma 3.4.
Suppose $\mathcal{A}$ is an Abelian category with $\mathrm{hd}(\mathcal{A})\in\{0,1\}$ and consider an admissible subcategory $\mathcal{B}\subset\mathrm{D}^{\mathrm{b}}(\mathcal{A})$. $\mathcal{A}\cap\mathcal{B}$ is the heart of a bounded t-structure on $\mathcal{B}$.
Proof.
Consider $X\in\mathcal{B}$ and write $X=\bigoplus_{k\in\mathbb{Z}}H^{k}(X)[-k]$ as in 3.3. $\mathcal{B}$ is closed under summands and shifts, so $H^{k}(X)\in\mathcal{B}$ for all $k$. On the other hand, associated to the standard t-structure is a filtration of $0=X_{n+1}\to X_{n}\to\cdots\to X_{2}\to X_{1}=X$ such that $\operatorname{Cone}(X_{i+1}\to X_{i})$ is given by one of the nonzero $H^{k_{i}}(X)[-k_{i}]$. It follows that each $X_{i}$ lies in $\mathcal{B}$ and consequently that $\mathcal{A}\cap\mathcal{B}$ is the heart of a bounded t-structure on $\mathcal{B}$. ∎
Proposition 3.5.
Suppose $\mathcal{A}$ and $\mathcal{B}$ are as in 3.4. There exists an exact equivalence $r:\mathrm{D}^{\mathrm{b}}(\mathcal{A}\cap\mathcal{B})\to\mathcal{B}$ which restricts to the identity on $\mathcal{A}\cap\mathcal{B}$.
Proof.
By [LPZ]*Lem. 3.1, there is a t-exact functor $r:\mathrm{D}^{\mathrm{b}}(\mathcal{A}\cap\mathcal{B})\to\mathcal{B}$ which restricts to the identity on the heart $\mathcal{A}\cap\mathcal{B}\subset\mathcal{A}$ of 3.4. By Lemma 3.3 ibid. this is an exact equivalence, since given a pair of objects $X,Y$ of $\mathcal{A}\cap\mathcal{B}$, one has $\operatorname{Hom}_{\mathrm{D}^{\mathrm{b}}(\mathcal{A})}(X,Y[n])=0$ for all $n\not\in\{0,1\}$. ∎
Theorem 3.6.
If $\mathcal{A}$ is a noncommutative curve, then any admissible subcategory of $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ is equivalent to the bounded derived category of a category of type (1) or (2) in 2.5. In particular, $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ contains no phantom subcategories.
Proof.
Let $\mathcal{B}$ be an admissible subcategory of $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ and consider $X,Y$ in the associated heart $\mathcal{A}\cap\mathcal{B}$ by 3.4. By 3.5 one sees $r(X)=X$ and $r(Y[n])=Y[n]$. Consequently, $\operatorname{Hom}_{\mathrm{D}^{\mathrm{b}}(\mathcal{A}\cap\mathcal{B})}(X,Y[n])=\operatorname{Hom}_{\mathrm{D}^{\mathrm{b}}(\mathcal{A})}(X,Y[n])$ for all $n\in\mathbb{Z}$. It follows that $\mathrm{hd}(\mathcal{A}\cap\mathcal{B})\in\{0,1\}$. $\mathcal{A}\cap\mathcal{B}$ is Noetherian, being a full Abelian subcategory of $\mathcal{A}$. Finally, saturatedness is a property inherited by semiorthogonal factors [B-KSerre]*Prop. 2.8 so by 3.5 we conclude that $\mathcal{A}\cap\mathcal{B}$ is of the claimed form.
It remains to observe that each of the categories listed in 2.5 has nontrivial $K_{0}$. This follows from the existence of $\dim:K_{0}(\operatorname{rep}Q)\to\mathbb{Z}$ and $\mathrm{rank}:K_{0}(\operatorname{Coh}(\mathcal{X}))\to\mathbb{Z}$, both of which are nonzero. Thus, $K_{0}(\mathcal{B})$ has positive rank and $\mathcal{B}$ is not a phantom category. ∎
Next, we consider the problem of classifying admissible subcategories of the derived categories of some noncommutative curves. When $X$ is a smooth projective curve over $k$ of genus at least one, a result of Okawa [Okawaindecomp] states that there are no nontrivial admissible subcategories of $\operatorname{D^{b}_{coh}}(X)$. On the other hand, the admissible subcategories of $\operatorname{D^{b}_{coh}}(\mathbb{P}^{1})$ are all of the form $\langle\mathcal{O}(n)\rangle$ for $n\in\mathbb{Z}$.
Corollary 3.7.
Let $Q$ denote a finite acyclic quiver. If $\mathcal{B}\subset\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q)$ is a nontrivial admissible subcategory then $\mathcal{B}\simeq\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q^{\prime})$ for $Q^{\prime}$ a finite acyclic quiver.
Proof.
By the proof of 3.6, $\mathcal{B}$ is equivalent as triangulated categories to $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q\cap\mathcal{B})$, where $\operatorname{rep}Q\cap\mathcal{B}$ is one of the categories classified in 2.5. As an Abelian subcategory of $\operatorname{rep}Q$, $\operatorname{rep}Q\cap\mathcal{B}$ is also of finite length and thus by 2.7 equivalent to $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q^{\prime})$ for $Q^{\prime}$ another finite acyclic quiver. ∎
If $Q$ is a finite acylic quiver with $n$ vertices, then the simple representations give a full exceptional collection of $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q)$. On the other hand, work of Crawley-Boevey [CrawleyBoevey] implies that the mutation action by $\mathbb{Z}^{n}\rtimes\mathrm{Br}_{n}$ on the full exceptional collections in $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q)$ is transitive. 3.7 implies that every admissible subcategory of $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q)$ admits a full exceptional collection and in particular we obtain a complete classification of the admissible subcategories of $\mathrm{D}^{\mathrm{b}}(\operatorname{rep}Q)$.
Given the above discussion, one might wonder if proper admissible subcategories of noncommutative curves always admit full exceptional collections. This is not the case as the following example shows:
Example 3.8.
Consider a degree $2$ branched cover $f:X\to Y$ where $X$ and $Y$ are smooth projective curves and $g(Y)\geq 1$. The Galois group of the covering is $G=\mathbb{Z}/2\mathbb{Z}$ and there is a commutative diagram
${X}$${{[}X/G{]}}$${Y}$$\scriptstyle{f}$$\scriptstyle{\pi}$
where $\pi$ is the coarse moduli space map. $[X/G]$ is an orbifold curve in the sense of 2.4. Using the fact that the category of $G$-equivariant coherent sheaves on $X$ is equivalent to $\operatorname{Coh}([X/G])$ combined with [CP10]*Thm. 5.1, one sees that $\pi^{*}$ identifies $\operatorname{D^{b}_{coh}}(Y)$ with an admissible subcategory of $\operatorname{D^{b}_{coh}}([X/G])$. Since $g(Y)\geq 1$, this is an indecomposable admissible subcategory by [Okawaindecomp].
4. Alternative proof of nonexistence of phantoms
In this last section, we present the original argument for the nonexistence of phantoms for noncommutative curves from an earlier version of this paper. The merit of this argument is that the proof is logically independent of [ReitenvDB] and thus perhaps more elementary.
Theorem 4.1.
If $\mathcal{A}$ is a noncommutative curve then $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ contains no phantom subcategories.
Definition 4.2.
An Abelian category $\mathcal{A}$ is called visible if for all $X\in\mathcal{A}$, $[X]=0\in K_{0}(\mathcal{A})$ implies $X=0$.
The following simple lemma is the main observation.
Lemma 4.3.
If $\mathcal{A}$ is a visible Abelian category with $\mathrm{hd}(\mathcal{A})\in\{0,1\}$, then $\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ contains no phantom subcategories.
Proof.
Suppose $\mathcal{B}\subset\mathrm{D}^{\mathrm{b}}(\mathcal{A})$ is an admissible subcategory. Let $X\in\mathcal{B}$ be given and decomposed as in 3.3. If $X\neq 0$, then there exists an $i$ such that $H^{i}(X)\neq 0$ and by 3.2 one has $\mathcal{A}\cap\mathcal{B}\neq 0$. If $\mathcal{B}$ were a phantom then its objects would have zero image in $K_{0}(\mathcal{A})$. In particular, any $X\in\mathcal{A}\cap\mathcal{B}$ must have trivial class in $K_{0}(\mathcal{A})$. By hypothesis, this implies $\mathcal{A}\cap\mathcal{B}=0$ and hence that $\mathcal{B}=0$. ∎
We next prove that examples of interest satisfy the visible condition.
Lemma 4.4.
If $\mathcal{A}$ is a finite length Abelian category then $\mathcal{A}$ is visible.
Proof.
$K_{0}(\mathcal{A})=\bigoplus_{S\>\text{simple}}\mathbb{Z}\cdot[S]$ and given $M\in\mathcal{A}$, $[M]=\sum_{i=1}^{k}n_{i}[S_{i}]$ where $S_{1},\ldots,S_{k}$ are the simples appearing in any composition series of $M$ with multiplicities $n_{1},\ldots,n_{k}$, respectively. Visibility follows. ∎
Corollary 4.5.
If $A$ is a hereditary finite dimensional algebra over a field $k$ then $\mathrm{D}^{\mathrm{b}}(\mathrm{mod}\>A)$ contains no phantom subcategory.
Proof.
$\mathrm{mod}\>A$ is finite length so we apply 4.4 and 4.3. ∎
We suppose for the next result that $k$ is algebraically closed and of characteristic zero.
Proposition 4.6.
If $\mathcal{X}$ is a projective Deligne-Mumford stack, then there exists a group homomorphism $P:K_{0}(\mathcal{X})\to\mathbb{Q}[t]$ such that for $\mathcal{F}\in\operatorname{Coh}(\mathcal{X})$, $P(\mathcal{F})=0$ implies $\mathcal{F}=0$.
Proof.
Let $\pi:\mathcal{X}\to X$ denote the coarse moduli space map of $\mathcal{X}$. For a locally free generating sheaf $\mathcal{E}$ on $\mathcal{X}$ (such a sheaf always exists by the projectivity assumption [Kresch]*Thm. 5.3), there is a functor $F_{\mathcal{E}}:\operatorname{QCoh}(\mathcal{X})\to\operatorname{QCoh}(X)$ given by $\mathcal{F}\mapsto\pi_{*}\mathcal{H}\mathrm{om}_{\mathcal{O}_{\mathcal{X}}}(\mathcal{E},\mathcal{F})$ [Nironi09]*Defn. 2.4. As $\pi$ is proper, $F_{\mathcal{E}}$ induces a functor $\operatorname{Coh}(\mathcal{X})\to\operatorname{Coh}(X)$.
$F_{\mathcal{E}}$ is exact and for all $\mathcal{F}\in\operatorname{Coh}(\mathcal{X})$, $F_{\mathcal{E}}(\mathcal{F})=0$ if and only if $\mathcal{F}=0$ [Nironi09]*Lem. 3.4. Fix a very ample line bundle $\mathcal{O}_{X}(1)$ on $X$. For $\mathcal{G}\in\operatorname{Coh}(X)$ let $\mathcal{G}(m):=\mathcal{G}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{X}(1)^{\otimes m}$. Using this, one defines $P_{\mathcal{E}}(\mathcal{F},m):=\chi(X,F_{\mathcal{E}}(\mathcal{F})(m))$, which is a rational polynomial function of $m\in\mathbb{Z}$ and defines a polynomial $P_{\mathcal{E}}(\mathcal{F},t)\in\mathbb{Q}[t]$. This gives a homomorphism $P_{\mathcal{E}}(-):K_{0}(\mathcal{X})\to\mathbb{Q}[t]$. Now, $P_{\mathcal{E}}(\mathcal{F},t)\in\mathbb{Q}[t]$ is the Hilbert polynomial of $F_{\mathcal{E}}(\mathcal{F})\in\operatorname{Coh}(X)$; so, if $P_{\mathcal{E}}(\mathcal{F},t)=0$, then $F_{\mathcal{E}}(\mathcal{F})=0$ and hence $\mathcal{F}=0$. ∎
Corollary 4.7.
If $\mathcal{X}$ is a projective orbifold curve, then $\operatorname{Coh}(\mathcal{X})$ is visible. Consequently, $\operatorname{D^{b}_{coh}}(\mathcal{X})$ contains no phantom subcategories.
Proof.
$\operatorname{Coh}(\mathcal{X})$ is visible by 4.6; indeed, for $\mathcal{F}\in\operatorname{Coh}(\mathcal{X})$, if $[\mathcal{F}]=0$, then $P_{\mathcal{E}}(\mathcal{F})=0$ and $\mathcal{F}=0$. We conclude by 4.3. ∎